13. Find the standard reduction potentials for the following half-reactions: So+4H + 2e-> HSO H2O and Ag + e- > Ag a) Write the balanced overall reaction for a successful cell made from these two couples. b) Write the line notation for the cell. c) What is Eo for the cell? d) What is the equilibrium constant for the cell reaction at 250C? e) Calculate the ratio of activities of produces and reactants, Q, that will produce a cell voltage of 0.51 V.
Answer #1

13. Electrochemical cell

a) balanced overall reaction for a sucessful cell made from these two couples

2Ag+ + H2SO3 + H2O ---> SO4^2- + 4H+ + 2Ag

b) line notation for the cell

Ag+(aq) | Ag(s) || H2SO3(aq) | SO4^2-(aq),H+(aq)

c) Eo = Ecathode - Eanode

         = 0.80 - 0.172

         = 0.628 V

d) equilibrium constant Kc at 25 oC

T = 25 oC + 273 = 298 K

R = gas constant

n = 2

F = Faraday constant

Using,

dGo = -nFEo = -RTlnKc

Kc = nFEo/RT

     = inv.ln(2 x 96500 x 0.628/8.314 x 298)

     = 1.76 x 10^21

e) when cell voltage E = 0.51 V

Using Nernst equation,

E = Eo - 0.0592/n log[activity of products/activity of reactants]

n = 2

feeding values,

[activity of products/activity of reactants] = inv.log[(0.628 - 0.51) x 2/0.0592]

                                                                 = 9694